Thank you very much for all you do to help those who struggle!
19. Adult tickets for a play cost $15 and child tickets cost $12. If there were 26 people at a performance and the theater collected $336 from ticket sales, how many adults and how many children attended the play?


Answer by stanbon(75887)

Adult tickets for a play cost $15 and child tickets cost $12. If there were 26 people at a performance and the theater collected $336 from ticket sales, how many adults and how many children attended the play?
---------------------
Let number of adult tickets be "x" ;Value is 15x dollars
Number of child tickets is "26-x"; Value is 12(26-x)=312-12x dollars
-------------
value + value = 336
15x+312-12x=336
2x=24
x=12 (number of adult tickets)
26-x=26-12=14 (number of child tickets)
Cheers,
Stan H.